OpenStudy (redneckoutlaw):

10.06] What is the product of the square root of 2. (6bthe square root of 16. − 2b)? Simplify if possible. Choose one answer. a. 16b b. 4bthe square root of 2. c. 22bthe square root of 2. d. 16bthe square root of 2.

OpenStudy (aum):

\[ \sqrt{2}*\left( 6b\sqrt{16} - 2b\right ) = \sqrt{2}*(6b * 4 - 2b) = \sqrt{2}*(24b - 2b) = \sqrt{2}*(22b) = 22b\sqrt{2} \]

OpenStudy (redneckoutlaw):

Mind helping mewith a few more?